BROADENING AND SHIFT COEFFICIENTS IN THE ν_{3} BANDS OF ^{12}C^{16}O_{2} AND ^{13}C^{16}O_{2}

\underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; D. CHRIS BENNER, V. MALATHY DEVI, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795.

In a previous study^{a} we had reported N_{2}-broadening and pressure-induced shift coefficients for 34 rovibrational transitions in the ^{12}C^{16}O_{2} ν_{3} fundamental band near 4.3 μm. These parameters were determined from spectra recorded with the McMath-Pierce Fourier transform spectrometer (FTS) of the National Solar Observatory on Kitt Peak, Arizona. We now report similar measurements of N_{2}-broadening and shifts for transitions up to J′′ = 56 in the ^{13}C^{16}O_{2} ν_{3} fundamental band, plus determinations of N_{2}-broadening coefficients in the ^{13}C^{16}O_{2} ν_{2} + ν_{3} − ν_{2} hot band and the ^{13}C^{16}O^{18}O ν_{3} fundamental band. We also made new measurements of N_{2}-broadening and pressure-induced shift coefficients for rovibrational transitions up to J′′ = 46 in the ^{12}C^{16}O_{2} ν_{3} fundamental band. These results were obtained from simultaneous analysis of five absorption spectra using a multispectrum nonlinear least-squares technique^{b}. A 4.08 cm sample cell at room temperature was used to record all of the spectra at 0.003 cm^{−1} resolution with the McMath-Pierce FTS. This data set includes one low pressure (0.15 Torr) spectrum obtained with a 90% ^{13}C-enriched CO_{2} sample and four spectra of lean mixtures of the same ^{13}CO_{2} sample in N_{2}. Total pressures of the mixtures were between 101 Torr and 464 Torr. Because of the isotopic sample used, the ν_{3} fundamental bands of ^{12}CO_{2} and ^{13}CO_{2} appeared together in the same spectra, and we were able to obtain a consistent set of line parameters for both molecules. The present measurements represent the first experimental determination of N_{2}-broadening and pressure-induced shift coefficients in isotopic bands of CO_{2} in the 4.3 μm region. The results obtained for the various bands will be compared with each other, with the values in the HITRAN database^{c}, and with available values reported in the literature.

aV. Malathy Devi, D. Chris Benner, C. P. Rinsland and M. A. H. Smith, JQSRT \underline{48}, 581-590 (1992).

bD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

cL. S. Rothman, R. L. Hawkins, R. B. Wattson and R. R. Gamache, JQSRT \underline{48}, 537-566 (1992).